Let  F be the distribution function defined by   F(x) = 0 if 1+r if 2+x? if 19 if I <-1 -1 <r <0 0<r<2 > 2

Let  \mu _{F} be the Lebesgue-Stieltjes measure asociated to F .

Determine the measurements  of the fpllowing sets:

\left \{ 2 \right \}

[-\frac{1}{2},3)

(-1, 0]υ (1,2)

10, ) U (1, 2

(1 : 2) + 2x² > 1)
Answer #1

Solution Given function is ro if x.4-1 F(1 = htxe if IERO : Tex² if osaa2 19 if x22 following Nao we have determine the sets- 1- (1) = 1-() 아을 - 905 ( (-1, 0] (1,2) ㅋ 4 ((-1,]) = F (0) F (1) = 2 0 22. LA (1,2)) = f (2) - F ) = (2+2) ~ (2+3) = (+y-2M(4) [otju (1,2] → (10, 1)) = f))-FC0) = (2+)-1 → u (12]) = F (21)-F (1) = 9-3 . u ([0, b) U(1,2]) = +6 = 5+2453 (x: +22 > When, so, > JU1호의 (E, JJ) = F (J) F(-) ⇒ 살 | ((, )) )) - F (c) - F( - -- 4 675St.9+5.0 = st:9+ = (loc. {]n (10-))